OpenStudy (linn99123):

What is the slope of the line that passes through the points (–2, 15) and (–8, –5)? A.) -10/3 B.) -5/2 C.) 5/2 D.)10/3

undeadknight26 (undeadknight26):

-5 - (-15)/-8 - (-2)

OpenStudy (anonymous):

Just remember the rule when you add two negatives together

OpenStudy (solomonzelman):

Why do you use the second point as (x1,y,2) instead of the first ;) Doesn't matter, just curious.

OpenStudy (solomonzelman):

I mean (x,1,y1)

OpenStudy (anonymous):

Yes, I have learned it y2-y1 / x2-x1

OpenStudy (anonymous):

(x1, x2), (y1), y2)

undeadknight26 (undeadknight26):

@SolomonZelman i get 10/-6...where did i go wrong mate?

OpenStudy (solomonzelman):

ggg \[15-(-5)=20~~~~~~~-5-15=-20\]

OpenStudy (anonymous):

how did you get -6?

OpenStudy (solomonzelman):

either way ...

OpenStudy (anonymous):

-8 - (-2)

OpenStudy (anonymous):

oh, I answered my own question

OpenStudy (solomonzelman):

yes, so it is ±20 going whichever way...

undeadknight26 (undeadknight26):

-8 - -2 = -6...and ohhh... i did -5 - -15 lolz

OpenStudy (solomonzelman):

yes, and that's where you got confused, because the 15 is positive in the coordinate.

OpenStudy (anonymous):

it is easy to make an error when you are doing the subtraction

OpenStudy (anonymous):

but I find if you write it out with the - in () then you can just change it a plus

OpenStudy (solomonzelman):

it is (-2,15) , NOT (-2,-15)

undeadknight26 (undeadknight26):

Yeah i saw that lolz...

OpenStudy (solomonzelman):

good :)

OpenStudy (anonymous):

is there a way to field questions into another subject? there is a lot of questions all over the place, biology, chemistry, etc...

OpenStudy (anonymous):

I feel like I am helping others in multiple subjects in a math community

OpenStudy (linn99123):

i have no clue wat yall just did i tryed to figure it out but still didnt get the answer

OpenStudy (linn99123):

Its B?

undeadknight26 (undeadknight26):

-5 - 15/ -2 - -8 -20/-6 = -10/3 I would say A am i right @SolomonZelman

OpenStudy (anonymous):

you would have -20/6 right?

OpenStudy (anonymous):

-2 - - 8 = -2 + 8 = 6

OpenStudy (solomonzelman):

Yes, undeadknight, it's A